# Enhancing Multimodal Medical Image Classification using Cross-Graph Modal Contrastive Learning

### v4

@@ V-B Compare with the SOTA Methods
GCN+GCN 2-layer CNN 0.81$\pm$0.01 0.60$\pm$0.06 0.73$\pm$0.08 0.74$\pm$0.20 0.73$\pm$0.01 0.80$\pm$0.04 0.81$\pm$0.03 0.92$\pm$0.02
GAT+GAT 0.79$\pm$0.02 0.46$\pm$0.13 0.60$\pm$0.12 0.69$\pm$0.20 0.71$\pm$0.03 0.69$\pm$0.09 0.80$\pm$0.03 0.93$\pm$0.01
GCN+GCN ResNet18 0.87$\pm$0.01 0.65$\pm$0.01 0.76$\pm$0.03 0.75$\pm$0.05 0.78$\pm$0.02 0.76$\pm$0.02 0.86$\pm$0.02 0.95$\pm$0.01
GAT+GAT 0.87$\pm$0.01 0.67$\pm$0.03 0.74$\pm$0.03 0.75$\pm$0.03 0.78$\pm$0.02 0.77$\pm$0.02 0.87$\pm$0.01 0.94$\pm$0.01
Denotes the average of accuracy.
VI Ablation Study
In this section, we analyzed the impact of various parameters on CGMCL, including the classification performance with different values of the contrastive loss parameter $\beta$ across multiple backbone combinations and the parameter $K$ used for constructing the multi-modality neighborhoods graph, as shown in Supplementary Fig. S1.
VI-A Parameters of $\beta$ in the Objective Function
From a two-modality fusion perspective, we explore the influence of graph-based contrastive loss on feature representations in two domains. In Eq. 22, we introduce a parameterized $\beta$ as a weighting coefficient for $\mathcal{L}_{m}+\mathcal{L}_{f}$ and $\mathcal{L}_{contrastive}$. As Fig. 3, various CNN-based models serve as backbones, integrated with either GCN or GAT for graph-structured feature learning. Our results indicate that the combinations of ResNet18 with GAT or GCN achieve an AUC nearing 0.90 across different $\beta$ values in the MA vs. Abnormal and Normal vs. Abnormal classifications. However, in the most challenging task, Normal vs. MA, ResNet18 combined with GCN achieves a notable AUC of 0.80 at a $\beta$ value of 0.65.
VI-B Multimodal Features Alignment in KL Divergence Converge
@@
TABLE VI: Evaluating classification performance and ablation study of module components in CGMCL
@@
VI-C Module Ablation
To validate CGMCL in module ablation for multimodal classification performance in latent space, we compare (1) w/o Concatenate, which excludes the concatenation from Eqs. 9, and  10, (2) w/o IMFES, which removes the IMFES module from Eqs. 11, and  12 under different backbones with the entire module. As shown in Table V, the experimental results indicate that the GAT+GAT combination with a ResNet18 backbone yields a slight AUC increase for the more distinct Normal vs. Abnormal subtypes, compared to models without concatenation or IMFES. Notably, CGMCL demonstrates a 5.5% improvement in the Normal vs. MA classification. Furthermore, it can inferred that utilizing various multi-level concatenation layers and weighted inter-modality distributions effectively enhances the extraction of critical feature information.
VI-D Objective Function Ablation
We propose that a well-designed contrastive loss within the cross-graph modal objective function can effectively integrate feature spaces during the overall model optimization process. To evaluate CGMCL’s multimodal fusion performance, we tested various combinations of loss functions. These include the binary loss for single graph module fusion ($\mathcal{L}_{I}$ and $\mathcal{L}_{C}$), the addition of contrastive loss ($\mathcal{L}_{contrastive}$), and the comprehensive weighted loss CGMCL ($\mathcal{L}_{CGMCL}$). The experimental results in Fig. 4 show that $\mathcal{L}_{CGMCL}$ outperforms the combination of $\mathcal{L}_{I}$ + $\mathcal{L}_{C}$ by 2.1% in AUC and 2.2% in ACC for the MA vs. Abnormal classification task. Similarly, for the Normal vs. Abnormal task, $\mathcal{L}_{CGMCL}$ shows improvements of 1.1% in both AUC and ACC. These results indicate that the similarity between the two modalities for these subtypes is nearing the model’s convergence range. More importantly, the ability to distinguish early-stage PD is critical. In the Normal vs. MA classification task, $\mathcal{L}_{CGMCL}$ surpasses the next best-performing unweighted model $\mathcal{L}_{CGMCL}$ by 5.1% in AUC and 4.3% in ACC. This substantial improvement highlights CGMCL’s superior capability in capturing latent patterns characteristic of early-stage PD within the fusion space.
@@ VI-D Objective Function Ablation
Figure 5: Visualization of CGMCL predictions incorporating six indicators from twelve DaTQUANT parameters.
Figure 6: t-SNE visualization of CGMCL low-dimensional embeddings on two multimodal datasets.
VII Visualization of Multimodal Representation
VII-A Identification of PD subtype trajectories
For multimodal feature fusion, we present visual insights into the embeddings from the cross-graph modal fusion of patient images and meta-features generated by CGMCL. The final layer outputs of CGMCL serve as the two-dimensional embeddings for t-SNE visualization as shown in Fig. 5 and Fig. 6. We then project the corresponding images onto the low-dimensional data points. Fig. 6 (A) illustrates the low-dimensional embedding trajectories based on three severity classifications. The results show that CGMCL distinguishes between normal and abnormal cases in this reduced space. Our primary focus is the Normal vs. MA distinction, which is critical for early prevention. Our low-dimensional embedding demonstrates improved precision in differentiating subtle abnormalities from normal cases, addressing potential errors in manual classification.
Figure 7: Age-related regression estimates of CGMCL predictions for three PD subtypes.
(a) Normal vs. MA
(b) MA vs. Abnormal
(c) Normal vs. Abnormal
@@ VII-A Identification of PD subtype trajectories
VII-B Multi-Class Trajectories in Melanoma Embedding
@@
For multi-class melanoma classification, the seven-point criteria categorize skin lesion characteristics into groups based on similarity, with physicians assigning corresponding scores. Fig. 6 (B) illustrates how CGMCL generates low-dimensional representations based on similar features for PIG, PN, and STR, which correspond to feature labels (ABS, RIG, IR), (ABS, TYP, ATP), and (ABS, REG, IR), respectively. In Fig. 6, for PN, the feature score of ATP is 2 (e.g., atypical pigment network, blue-whitish veil, and irregular vascular structures). It is evident that ATP has a more distinct and severe scoring, showing clear separation from ABS and TYP. Some ATP positions are located within TYP, indicating that certain TYP cases may progress toward more malignant features. Compared to low-dimensional representations in previous studies, our CGMCL model integrates more precise multimodal information into feature representations and provides diagnostic insights, addressing the lack of interpretability often associated with meta-features in existing research work.
VIII Discussion
VIII-A Clustering Results
In this section, we assess the integration of diverse domains by evaluating CGMCL’s clustering capabilities for three subtypes of PD in similarity representations across various patient modalities. We visualize the patient similarity matrix derived from Eq. 13 for PD and melanoma in Fig. 9. Notably, CGMCL achieves near-perfect clustering between Normal and Abnormal cases in Fig. 9 (a). However, the clustering boundaries become less distinct when differentiating between Normal vs. MA and MA vs. Abnormal (e.g., prodromal and moderate abnormal PD). The broader clustering boundary range for Normal vs. MA in the prodromal stage indicates that CGMCL can, to some extent, distinguish prodromal PD patients. In the melanoma clustering analysis, as shown in Fig. 9 (b), the classification of different classes reveals three distinct community clusters in the similarity matrices of DaG and PIG across all categories. This demonstrates CGMCL’s ability to differentiate feature characteristics in multimodal melanoma, even when dealing with samples that have fewer classes.
VIII-B Neurodegenerative Subtype Analysis in PD
Neurodegenerative conditions have long posed a predictive challenge in PD. Our proposed CGMCL aims to predict better and analyze the age-related effects on neurodegenerative PD subtypes. The scatter plot and regression estimates in Fig. 7 (left) reveal a weak but statistically significant correlation ($p<0.05$) between age and the model’s predictions for distinguishing Normal from MA. A subtle upward trend indicates that the model tends to classify older individuals as MA with a slightly higher probability. In contrast, the analysis of MA vs. Abnormal in Fig. 7 (middle) shows that age accounts for approximately 33.1% of the variance in model predictions, a finding that is highly statistically significant ($p<0.001$). This result indicates that age decreases the likelihood of being predicted as fully abnormal rather than mildly abnormal. Furthermore, Fig. 7 (right) demonstrates that the model’s predictions between Normal and Abnormal subtypes are not statistically significant ($p>0.05$), indicating that age is not a key factor in distinguishing these two categories. Although there is a slight increase in the probability of being classified as Abnormal with advancing age, the likelihood of being categorized as fully abnormal decreases with age. Based on these observations, the impact of age on differentiating abnormal cases appears minimal in late-stage PD, both in terms of manual interpretation and imaging features.
VIII-C Explainable Modality Diagnosis
Traditional SPECT imaging for PD relies on manual interpretation and diagnosis by nuclear medicine physicians. In comparison to commonly used Grad-CAM methods (e.g., CNN visualization) for diagnosis [37], we developed a quantifiable diagnostic score that accurately predicts the magnitude of meta-features for patients at different stages. This score integrates our model’s predicted modality contribution with twelve meta-features, utilizing a scaled modality mechanism to calculate specific contribution scores [38]. In Fig. 8, we randomly selected patient samples from each subtype for analysis. We calculated the modality weight in CGMCL for the twelve meta-features, using importance scores given by $Score=X^{C}\odot\mathrm{IMFES}(\mathcal{H}^{C})$. Observing the Normal vs. MA comparison in Fig. 8 (a), patients beginning to exhibit MA characteristics (third image in Fig. 8 (a)) show a notably asymmetric geometric pattern in the striatal regions: PC-R, PC-L, and PP-L. An intriguing observation emerges when comparing the radar charts of normal patients with those of MA and Abnormal patients. Both groups with anomalies display substantially irregular radar chart shapes. This irregularity is particularly pronounced in the Abnormal group, where the overall area covered by the radar chart is smaller than in normal cases. This phenomenon is attributed to early-stage PD in patients, characterized by evident striatal atrophy in both hemispheres.
IX Conclusion
This study introduces a novel multimodal fusion framework that leverages cross-graph modal and integrates concatenated multi-level inner feature maps. Our approach effectively combines medical imaging data with clinical features, enhancing multimodal fusion. Additionally, we employ contrastive learning within the common latent space of fused same-modality data to improve the model’s classification accuracy for various subtypes across two multimodal datasets. Regarding clinical multimodal interpretability, our proposed CGMCL differs from existing methods by incorporating non-image features, particularly enabling quantitative clinical interpretation of SPECT imaging in early-stage PD. Furthermore, CGMCL shows potential for broader application in diverse multimodal disease studies.

### v5

@@ 4 Dataset Collection
[…]
Comparison with Unimodal Improvements (%) - +3.57%∗∗∗ +39.58%∗∗∗ +26.67%∗∗∗ +15.38%∗∗∗ +5.41%∗∗∗ +15.94%∗∗∗ +6.10%∗∗∗ +26.67%∗∗∗
Comparison with Multimodal Fusion Improvements (%) - +0.00% +1.52% +5.56%∗∗∗ +5.63%∗∗∗ +0.00% +5.26%∗∗∗ +2.35%∗∗∗ +25.00%∗∗∗
Comparison with Vision-Transformer Improvements (%) - +1.16%∗∗ +1.52% +2.70%∗∗∗ +0.00% +1.30%∗∗ +2.56% +0.00% +1.06%∗∗∗
Significance. p << 0.05, ∗∗p << 0.01, ∗∗∗p << 0.001 indicate statistically significant improvements of CGMCL compared to the corresponding baseline models. Denotes the average of accuracy.
Figure 5: Robustness evaluation of the proposed CGMCL model under Gaussian noise perturbations ($\sigma=0.01$$0.10$) for three SPECT-based PD subtyping tasks.
5 EXPERIMENTS
5.1 Experiments Setting
In this work, the experiments were conducted using the PyTorch framework on a workstation equipped with an NVIDIA GeForce RTX 3070 GPU (8 GB VRAM) with CUDA 12.2 support. The models were jointly optimized with the Adam optimizer (learning rate = 0.001) for 300 epochs and a batch size of 32. As shown in Table 3, CGMCL achieves the best balance between accuracy and computational efficiency. It has fewer parameters (15.07 M) and lower latency (8.879 ms) than other graph-transformer models, with comparable GFLOPs (68.052). Its training time (32.94 s) is shorter than most transformer-based baselines, demonstrating faster convergence. Although GPU usage (1997.5 MB) is slightly higher than GradFormer, it remains efficient relative to GTC and GEAET. Overall, CGMCL achieves high performance with low computational cost, indicating that the proposed cross-graph contrastive learning framework offers a good trade-off between accuracy and efficiency for multimodal medical imaging tasks.
5.2 Baseline Methods Comparison
To fairly compare the effectiveness of CGMCL across various methods for both the PD and melanoma datasets, we conducted a quantitative analysis of baseline method comparisons. The results for PD, segmented into three subtypes, are presented in Table 4 through Table 6, and for the melanoma multi-class dataset in Table 7. Initially, we employed conventional machine learning algorithms to assess the impact of a single modality on meta-features. This analysis applied logistic regression, XGBoost, random forest, support vector machines (SVM) Hsu_2019, and AdaBoost freund1996experiments to both datasets. To further evaluate the performance of our proposed CGMCL model in PD classification, we compared it to unimodal two-layer CNN models. We then extended the analysis using SPECT images with three slices to compare different 3D-CNN architectures. When comparing machine learning methods that utilize only meta-features against CNN-based unimodal approaches, nonlinear logistic and tree-based methods (i.e., XGBoost and AdaBoost) achieved accuracies between 0.58 and 0.63 for distinguishing Normal from MA in PD. This indicates that relying solely on quantitative data may result in losing other informative features, such as those derived from images. In contrast, unimodal image feature extraction models substantially improved classification accuracy for the three PD subtypes. Specifically, ResNet18 increased accuracy by 16% for Normal vs. MA, 12.5% for Normal vs. Abnormal, and 1.16% for Normal vs. Abnormal compared to logistic regression. Conversely, for the melanoma dataset, meta-features in ML methods outperform in the DaG, PN, and VS categories.
5.3 Compare with the SOTA Methods
In this work, we examined the effectiveness of our proposed CGMCL framework by conducting extensive comparisons with a range of state-of-the-art (SOTA) unimodal and multimodal models, such as MHCA, DeAF, and TriFormer. These models employ 3D tensor-based CNN architectures as feature extractors jointly trained with clinical meta-features, transformer-based multimodal fusion methods, and recent graph-transformer hybrid models. The quantitative results are presented for both the Parkinson’s disease dataset (Tables 4, 5, 6) and the melanoma dataset (Table 7). Our experimental results yielded the following findings:
  1. 1.
  2. 2.
    We further evaluated the cross-modal graph fusion capability of our proposed CGMCL model in classifying the three PD subtypes. From Table 4, using GAT as the cross-modal fusion method with ResNet18 achieves metric values between 0.73 and 0.76, comparable to the SOTA model using 3D-CNNs or unimodal approaches (ACC $\uparrow$ 4.1%, PPV $\uparrow$ 3.9%). This indicates that the fusion of clinical and imaging data can more effectively assist in identifying difficult-to-diagnose early PD symptoms. Moreover, there is a noticeable improvement in classification performance for the other two subtypes (e.g., MA vs. Abnormal and Normal vs. Abnormal).
  3. 3.
  4. 4.
    We compared CGMCL with recent graph-transformer architectures (GEAET, CoBFormer, GTC, GradFormer, Cluster-GT) that exhibit strong capability in learning structural dependencies among patients and vision-transformer models (SHViT, MedFormer, BIIGMA, PLRR). As shown in Tables 5 to 7, CGMCL achieves competitive performance across PD subtypes, with notable improvements in Normal vs. MA classification (ACC=0.76, +2.70% over best graph-transformer baseline). For melanoma (Table 7), CGMCL demonstrates superior multi-class classification in five categories (DaG, PIG, PN, VS, DIAG), achieving 95% ACC in the five-class DIAG task versus PLRR (94%) and BIIGMA (92%).
5.4 Robustness Analysis
Unlike the popular high-quality publicly available PPMI dataset for PD analysis, prior research on PPMI data has shown that high-quality data often leads to improved model accuracy and stability ding2021dopamine. However, there is a lack of more refined early-stage PD diagnosis (e.g., MA) and realistic noisy medical images commonly found in most hospitals. Therefore, to simulate common image degradation scenarios in clinical practice, we added varying levels of Gaussian noise ($\sigma=0.00$-$0.10$) to the input images to analyze the classification stability of CGMCL, as shown in Fig. 5. Compared to baseline models (GEAET, CoBFormer, GTC, GradFormer, Cluster-GT-T), which exhibited significant performance degradation under the same noise conditions with accuracy drops of 10–20%, CGMCL demonstrated superior robustness. When noise intensity increased from $\sigma=0.00$ to $\sigma=0.10$, CGMCL’s accuracy only fluctuated slightly from 0.75 to the 0.67–0.77 range in the Normal vs. MA task, maintained a high level of 0.84–0.88 in the MA vs. Abnormal task, and remained stable at 0.87–0.90 in the Normal vs. Abnormal task. This robustness stems from the cross-graph contrastive learning and inter-modality feature scaling, which jointly enhance noise tolerance and feature consistency.
6 Ablation Study
To evaluate the contribution of each component to CGMCL, we conduct ablation studies examining the effectiveness of different components on overall classification performance. Specifically, we analyze: (1) the alignment of multimodal representations (Section 6.1), (2) the sensitivity of the parameter $K$ for $K$-nearest neighbors used in constructing the modality graph (Section 6.2), and (3) the IMFES module and the influence of the balancing weight coefficient $\beta$ between supervised classification loss and cross-graph contrastive loss (Sections 6.3 and 6.4).
6.1 Multimodal Features Alignment in KL Divergence Converge
@@
Figure 7: The mean and standard deviation of our proposed CGMCL’s performance in terms of sensitivity and specificity for PD, based on varying numbers of K-neighborhoods for constructing the modality graph.
6.2 Impact of K-Neighborhood Graph Construction
Furthermore, to examine the robustness of our CGMCL model under different graph construction settings, we analyzed the influence of varying the number of $K$-neighborhoods in building the modality graph. As shown in Fig. 7, the mean and standard deviation of sensitivity and specificity were computed across different $K$ values for PD classification. The results indicate that the model maintains stable performance within a moderate range of $K$ (from 60 to 120), suggesting that CGMCL is relatively insensitive to small fluctuations in neighborhood size and effectively captures patient similarity relationships across modalities.
6.3 IMFES Module Ablation
To validate CGMCL in module ablation for multimodal classification performance in latent space, we compare (1) w/o Concatenate, which excludes the concatenation from Eqs. 9, and  10, (2) w/o IMFES, which removes the IMFES module from Eqs. 11, and  12 under different backbones with the entire module. As shown in Table 8, the experimental results indicate that the GAT+GAT combination with a ResNet18 backbone yields a slight AUC increase for the more distinct Normal vs. Abnormal subtypes, compared to models without concatenation or IMFES. Notably, CGMCL demonstrates a 5.5% improvement in the Normal vs. MA classification. Furthermore, it can inferred that utilizing various multi-level concatenation layers and weighted inter-modality distributions effectively enhances the extraction of critical feature information.
Table 8: Evaluating classification performance and ablation study of module components in CGMCL
@@ 6.3 IMFES Module Ablation
Figure 9: Visualization of CGMCL multimodal representations and clinical parameter analysis. (Left) Visualization of CGMCL predictions incorporating twelve indicators from twelve DaTQUANT parameters. (Right) t-SNE visualization of CGMCL low-dimensional embeddings on two multimodal datasets.
Figure 10: Grad-CAM visualization of multimodal feature maps across disease domains. Compared with unimodal and graph-transformer models, CGMCL produces more localized and clinically relevant Grad-CAM activations within striatal and lesion regions, indicating enhanced interpretability and cross-modal feature alignment.
6.4 Objective Function Ablation
The well-structured contrastive term within the cross-graph modal objective function plays a pivotal role in aligning multimodal feature distributions and enhancing fusion consistency during model optimization. We evaluate the combination of the supervised classification losses and the cross-graph contrastive loss with balancing parameter $\beta$ for CGMCL’s multimodal fusion performance.
(a) Effect of Loss Components
We first compared several configurations of the objective function to examine the effect of the contrastive and structural consistency losses. As illustrated in  Fig. 8, incorporating the contrastive loss significantly improved the alignment of multimodal representations. The results demonstrate that the full $\mathcal{L}_{CGMCL}$ yields the highest fusion performance. Specifically, for the MA vs. Abnormal classification, $\mathcal{L}_{CGMCL}$ surpasses $\mathcal{L}_{I}+\mathcal{L}_{C}$ by 2.1% in AUC and 2.2% in ACC, while for the Normal vs. Abnormal task, improvements of 1.1% are observed in both metrics. Notably, in the most challenging Normal vs. MA classification, which is critical for early Parkinson’s detection, the full weighted loss achieves a 5.1% AUC and 4.3% ACC gain over the next best unweighted setting. These improvements indicate that the contrastive component enhances the CGMCL’s ability to capture subtle multimodal correlations and latent disease patterns in the shared embedding space.
(b) Sensitivity of the Weighting Parameter $\beta$
We further investigated how varying $\beta$ affects the interplay between supervised and contrastive objectives. As illustrated in Fig. 6, models employing different CNN backbones (e.g., ResNet18, VGG19) in conjunction with GCN or GAT graph modules exhibit stable performance across a wide $\beta$ range. In the MA vs. Abnormal and Normal vs. Abnormal tasks, ResNet18 coupled with either GAT or GCN achieves an AUC close to 0.90 across $\beta\in[0.4,0.8]$, indicating robustness to weighting variations. For the more difficult Normal vs. MA task, optimal performance (AUC = 0.80) occurs at $\beta=0.60$, suggesting that moderate weighting of the contrastive term yields the best balance between modality discrimination and shared-space alignment. These ablation studies validate the necessity of an adaptive $\beta$ setting to ensure optimal cross-modal optimization and highlight the role of $\beta$ in maintaining equilibrium between intra-modal classification and inter-modal consistency.
7 Visualization of Multimodal Representation
7.1 Identification of PD subtype trajectories
As illustrated in Fig. 9 (Left), the embedding trajectory presents a smooth transition from Normal to MA and Abnormal, reflecting the clinical progression of dopaminergic degeneration. Notably, the CGMCL performance on the Normal versus MA distinction becomes a critical boundary for early disease prevention, suggesting improved sensitivity to subtle early-stage abnormalities. This enhanced separability indicates that the integration of imaging and quantitative DaTQUANT features effectively reduces representation ambiguity and supports early PD diagnosis. Furthermore, the final-layer representations of each patient, derived from the integrated image and meta-features, were embedded into a two-dimensional latent space as shown in Fig. 9 (Right). A particularly meaningful observation lies in the boundary region between Normal and MA clusters, which represents early-stage PD patients with subtle dopaminergic degeneration. In this transitional zone, the embeddings are less dispersed compared to unimodal baselines, suggesting that CGMCL successfully reduces representation ambiguity by leveraging complementary information from both SPECT images and DaTQUANT-derived quantitative metrics.
@@ 7.1 Identification of PD subtype trajectories
7.2 Multi-Class Trajectories in Melanoma Embedding
@@
For multi-class melanoma classification, the seven-point criteria categorize skin lesion characteristics into groups based on similarity, with physicians assigning corresponding scores. Fig. 9 (B) illustrates how CGMCL generates low-dimensional representations based on similar features for PIG, PN, and STR, which correspond to feature labels (ABS, REG, IR), (ABS, TYP, ATP), and (ABS, REG, IR), respectively. In Fig. 9, for PN, the feature score of ATP is 2 (e.g., atypical pigment network, blue-whitish veil, and irregular vascular structures). It is evident that ATP has a more distinct and severe scoring, showing clear separation from ABS and TYP. Some ATP positions are located within TYP, indicating that certain TYP cases may progress toward more malignant features. Compared to low-dimensional representations in previous studies, our CGMCL model integrates more precise multimodal information into feature representations and provides diagnostic insights, addressing the lack of interpretability often associated with meta-features in existing research work.
7.3 Heatmaps Visualization of Multimodal Feature Activations
To further evaluate the interpretability and feature localization capability of the proposed CGMCL framework across different disease domains. We use the Grad-CAM to visualize activation maps across disease domains, as shown in Fig. 10. For PD classification, CGMCL produces highly localized activations within the striatum, particularly in the caudate and putamen regions where dopaminergic degeneration occurs lopez2023dopaminergic, nazari2022explainable, saeed2020neuroimaging. In contrast, unimodal ResNet18 exhibits diffuse activations beyond striatal regions, while graph-transformer models show improved but inconsistent localization with scattered attention across multiple brain regions. Notably, in the normal vs. MA classification, CGMCL generates focused activations in the posterior putamen, consistent with early PD pathology. We can observe in melanoma dataset, the CGMCL significance captures lesion-specific morphological features, including concentrated activations on irregular pigmented structures for DaG, discriminative patterns for multi-class DIAG, and radial streaming patterns for STR. Fig. 10 shows that CGMCL’s unified multi-modality fusion yields more focused attention on disease-concentrated regions compared to other approaches.
8 Discussion
Multimodal medical image fusion faces challenges from heterogeneous distributions, prevalent noise, and the need for effective fusion methods for accurate diagnosis. Our proposed CGMCL framework integrates multimodal medical data through cross-graph modal fusion azam2022review, kalamkar2023multimodal. Specifically, CGMCL enables (1) cross-modal patient similarity clustering, and (2) quantitative meta-feature importance estimation for clinical decision support, and (3) anatomically interpretable attention visualization that aligns with established disease progression patterns.
As shown in Fig. 11, the learned similarity matrices capture distinct clustering patterns that quantify patient relationships in the latent space. For PD subtypes (Fig. 11 (a)), Normal vs. Abnormal displays strong diagonal concentration (similarity > 0.8), indicating well-separated embeddings, while Normal vs. MA shows more diffused patterns (0.4–0.6), consistent with the 76% vs. 90% accuracy gap in Tables 4 to 6 and highlighting the challenge of early-stage detection. The MA vs. Abnormal matrix exhibits intermediate similarity (0.7), suggesting gradual feature separation with disease progression. For melanoma (Fig. 11 (b)), despite class imbalance (RS and DaG having  30% fewer samples), CGMCL maintains concentrated diagonal patterns (> 0.75), demonstrating effective neighborhood aggregation and robustness across multimodal contexts.
To identify the neurodegenerative regions from PD, the Grad-CAM visualizations (Fig. 10) under four conditions (Normal, Normal vs. MA, MA vs. Abnormal, and Abnormal) consistently highlight the striatum, especially the posterior putamen drori2025multiparametric, ortiz2019parkinson. As disease severity increases, the activation extends anteriorly toward the caudate nucleus. This spatial pattern is consistent with established DAT imaging findings in PD, where dopaminergic loss begins in the posterior putamen and progresses anteriorly, typically showing lateral asymmetry lopez2023dopaminergic, calle2019identification, ogawa2018role, fiorenzato2021asymmetric.
Furthermore, CGMCL provides an effective visualization tool for meta-feature importance to support clinical decision-making, as shown in Fig. 12. For Normal vs. MA (Fig. 12 (a)), posterior putamen parameters (PP-R, PP-L) show highest importance (0.85-0.90), exceeding anterior putamen (0.60-0.65) and caudate (0.45-0.50), consistent with Lewy body pathology progressing from posterior to anterior regions lee2023substantia. For MA vs. Abnormal (Fig. 12 (b)), importance becomes balanced across regions (0.75-0.88), with elevated P/C ratio (0.82-0.85) and caudate parameters (0.80-0.83), reflecting widespread bilateral degeneration in advanced disease shin2007use, ciesielska2017depletion. For Normal vs. Abnormal (Fig. 12 (c)), all parameters show uniformly high importance (0.82-0.92), with striatal SBR reaching peak scores (0.91-0.92). This progression pattern demonstrates posterior putamen dominance in early detection, expanding to bilateral putamen-caudate involvement, and then global striatal changes pitcher2012reduced.
These experimental findings collectively confirm the effectiveness of our cross-graph modal fusion framework, which preserves intra-modal patient structures while achieving cross-modal alignment through contrastive learning, reducing KL divergence from 1.2 to 0.4–0.6 (Fig. 6 (b)). The IMFES module adaptively balances modality contributions, improving performance by 5.5% over simple concatenation (Table 8), while graph attention jointly models imaging and clinical similarities, enhancing reliability in the challenging Normal to MA gray zone. The Grad-CAM visualizations in Fig. 10 reveal that CGMCL yields more clinically meaningful attention patterns compared to single-modal baselines in the PD dataset. Specifically, the model’s attention shifts from bilaterally symmetric putamen in healthy controls to asymmetrically distributed striatal regions in PD patients, consistent with the characteristic dopaminergic degeneration patterns illustrated in Fig. 12.
9 Limitations and Future Work
Although the proposed CGMCL framework demonstrates promising performance in multimodal PD classification, certain limitations remain. In early-stage PD, patients typically exhibit asymmetric degeneration of the striatum, where one side tends to deteriorate first, but the affected side varies among individuals. Eventually, both sides decline to a similar level.
The case study results in Fig. 13 reveal several limitations that warrant further discussion. CGMCL performs inconsistently on MA cases representing early-stage PD. As shown in the left case of Fig. 13, the model misclassified an MA scan as Normal (false negative). The corresponding attention map indicates that the network primarily focused on high-intensity regions of the striatum while overlooking the subtle reduction in the putamen signal. Conversely, the right case illustrates the opposite error (over-sensitivity). Although the ground truth label was MA, the model predicted Abnormal (false positive), overemphasizing symmetric bilateral signal decreases. These findings suggest that the image modality tends to dominate during multimodal fusion, leading to an overreliance on pixel-level variations.
Figure 13: Case study on failure analysis in mildly abnormal PD classification with the proposed CGMCL model
In future work, we will address these issues by (1) incorporating region-specific graph nodes (e.g., putamen and caudate ROIs) to enhance spatial representation, and (2) developing a larger multimodal framework to improve fine-grained ROI detection for more precise classification. Alternatively, we will explore CLIP-based multimodal models radford2021learning, zhang2023biomedclip with clinical text data to further enhance interpretability and diagnosis.
10 Conclusion
This study introduces a novel multimodal fusion framework that effectively integrates medical imaging and clinical data through dual-graph modeling and contrastive learning within a unified latent space. Our approach effectively combines medical imaging data with clinical features, enhancing multimodal fusion. By aligning heterogeneous modalities in a unified latent space, CGMCL enhances feature representation, interpretability, and robustness in the classification of two multimodal datasets. Additionally, the IMFES module further mitigates modality imbalance, enabling the model to adaptively balance contributions from diverse data sources. These results show that CGMCL has strong potential for broader applications in diverse multimodal disease studies and future clinical decision-support systems.
CRediT authorship contribution statement
Jun-En Ding: Conceptualization, Methodology, Software, Writing – original draft, Writing – review & editing. Chien-Chin Hsu: Data collection and medical guidance. Chi-Hsiang Chu: Statistical guidance. Shuqiang Wang: Methodological guidance. Feng Liu: Supervision, Methodology, Validation.
10.1 Declarations
Ethics approval and consent to participate
This study was approved by the Institutional Review Board of Taipei Veterans General Hospital, Taipei, Taiwan. Given the retrospective nature of the study, the requirement for informed consent was waived by the IRB.
11 Acknowledgment:
We are grateful to the Department of Nuclear Medicine at Kaohsiung Chang Gung Memorial Hospital for providing us with comprehensive data and data labeling support.
11.1 Funding
This study received no funding support.